Consider a political discussion group consists of five democrats and six republicans. Four people are selected (at random) to attend a conference. In how many ways can four republicans be selected from the group?


1
Expert's answer
2021-07-06T16:30:19-0400

Total number of republicans N=6

n= 4

Number of ways "= \frac{N!}{n!(N-n)!}"

"= \frac{6!}{4!(6-4)!} = \frac{5 \times 6}{2}= 15"